find the dimension of a rectangle whose width is 2 inches less than half the length and whose area is 160 square inches

Answer by LinnW(1048)

set W = the width
L = the length
W = (1/2)L - 2
L*W = 160 or
L*((1/2)L - 2) = 160
(1/2)L^2 -2L = 160
multiply each side by 2
L^2 -4L = 320
L^2 -4L -320 = 0
(L-20)(L+16) = 0
L = 20 is the only positive answer.
Since W = (1/2)L - 2
W = 10 - 2
W = 8